# Bankrupting DoS Attackers

### v4

@@ Appendix 0.A Analysis of LINEAR
By the algorithm, $\hat{g}({\mathcal{I}_{j}})\geq 1$ for all $j$. Let $\mathcal{I}_{tot}=\cup_{1\leq j\leq\ell}\mathcal{I}_{j}$. Using the additive property of $\hat{g}$, we get $\hat{g}(\mathcal{I}_{tot})=\sum_{1\leq j\leq\ell}\hat{g}(\mathcal{I}_{j})\geq\ell$. But by our definition of $\gamma$, we know that $\hat{g}(\mathcal{I}_{tot})\leq\gamma(g(\mathcal{I}_{tot})+1)=\gamma(g+1)$. Thus, $\ell\leq\gamma(g+1)$.
Lemma 2
@@
Putting together equations 1 and 2 we get that $g(\mathcal{I})\leq\gamma+\gamma^{2}+1$.
Lemma 3
@@
which proves the lemma.
@@
[…]
@@
The second step holds since $(b/\ell)-1=(b-\ell)/\ell$. The fourth step holds since, by Lemma 1, $\ell\leq\gamma(g+1)$.
@@
Let $f:\mathbb{R}\rightarrow\mathbb{R}$ such that, $f(x)=\sum_{j=1}^{\beta}(x+j)^{\alpha}$. For any $i\in[1,\ell]$, this function is continuous when $x\in[b_{i},b_{i}+1]$ and differentiable when $x\in(b_{i},b_{i}+1)$, since it is the sum of a finite number of continuous and differentiable functions respectively. For any $i\in[1,\ell]$, by applying the Mean Value Theorem for the interval $[b_{i},b_{i}+1]$, we have:
[…]
@@
In the above, the last line follows by Lemma 3, and noting that $1/\alpha$ is a fixed constant.
@@
which completes the proof.
@@
Proof
Since there is no latency, each client immediately receives the current value of PRICE if its first message to the server has a fee with value too small. Then, on the second attempt the client attaches a fee with the correct value, and so receives service. Hence, each good job is serviced with at most $2$ messages from client to server and at most $1$ message from server to client.
Theorem 1.1 now follows directly from Lemmas 6 and 7.
Appendix 0.B Analysis of LINEAR-POWER
@@
Putting these together, we get that when $B_{i}=0$, $x\leq\log 3\gamma^{2}$ and when $B_{i}>0$, $x\leq((1/2)\log B_{i})+1+\log 3\gamma^{2}$. Hence the maximum current price is at most $3\gamma^{2}\max(1,2\sqrt{B_{i}})\leq 6\gamma^{2}\sqrt{B_{i}+1}$.
Lemma 9
@@
By Lemma 8, $2^{x}\leq 6\gamma^{2}\sqrt{B_{i}+1}$; taking the log and simplifying, we get $x\leq 3+2\log(\gamma(B_{i}+1))$. The lemma follows since over a period of $2\Delta$ seconds, at most $2M$ good jobs can arrive.
Lemma 10
@@
If neither the current epoch nor the preceding epoch spanned the beginning of an iteration, then the current price never resets to $1$ during the lifetime of any of these jobs. So the current price is non-decreasing during each such job’s lifetime. Thus, none of the jobs can ever have an attached fee that is larger than the current price, and so no overpayment occurs.
Lemma 11
@@
By Lemma 8, the maximum current price in either epoch $i$ or $i-1$ is:
@@
This is the maximum amount that any job serviced in epoch $i$ can overpay. Multiplying by the number of good jobs serviced, we obtain the bound in the lemma statement.
Lemma 12
@@
Lemmas 1314, and 15 below complete the proof of our second main result, Theorem 1.2.
Lemma 13
@@
Proof
Fix the order in which jobs are serviced by LINEAR-POWER. The current price when a job is serviced in LINEAR-POWER will differ by at most a factor of $2$ from LINEAR because of the fact that LINEAR-POWER uses powers of $2$. Thus, the total amount spent by the adversary decreases by at most a factor of $2$, and the total amount spent by the algorithm on all jobs that do not overpay increases by at most a factor of $2$. It follows that the asymptotic results in Theorem 1.1 hold, except for the costs due to overpayment by the good jobs.
So, we can bound the asymptotic cost of the algorithm as the cost from LINEAR plus the cost due to overpayment. The cost in the theorem statement is the asymptotic sum of the costs from Theorem 1.1 and from Lemma 12. This is:
@@
which completes the argument.
Lemma 14
@@
Every time the job is bounced, the latency increases by at most $2\Delta$ seconds: $\Delta$ seconds to send the job to the server, and $\Delta$ seconds to receive the new current price from the server. When the job is serviced it takes at most $\Delta$ seconds: the time to send the job to the server. Thus, the total time for the job to be serviced is at most $2\Delta(2\log(\gamma(B+1))+2)+\Delta\leq\Delta(4\log(\gamma(B+1))+4)$.
@@
Every time a client sends a new message to the server, the fee attached to that message doubles. The maximum current price set by the server over all iterations is at most $\log(B+3\gamma^{2})$, so a good job will send at most that number of messages. Hence, all $g$ good jobs send a total of $O(g\log(B+3\gamma^{2}))$ messages to the server.
Appendix 0.C Lower Bound Proof
@@ 0.C.2 Analysis
$=\hat{g}(\mathcal{I})$
In the above, the first line follows from Equation 6, and the last line follows from the fact that $\gamma\geq 1$, and the definition of the estimator. Hence, by the above analysis, our estimator has $\gamma$ estimation gap for both distributions.
@@
To show $E(A-\sqrt{Bg\gamma})\geq 0$, we will show that $E(A)/\sqrt{E(B)g}\geq\sqrt{\gamma}$, and then cross-multiply and use Jensen’s inequality. If we define a random variable $a_{i}$ that has value $c_{i}$ if the $i$-th job is good and 00 otherwise, and note that $A=\sum_{i=1}^{n}a_{i}$, then by linearity of expectation, the expected algorithm cost due to fees in distribution $1$ is $\sum_{i=1}^{n}E(a_{i})=C\gamma g_{0}/n$, with service cost of $n$. A similar analysis for $E_{1}(B)$ gives:
@@
where the second line holds by applying Jensen’s inequality [89] to the convex function [^1] Equivalently, we can directly apply the (less well-known) version of Jensen’s inequality for concave functions [42] to obtain the second line. $f(x)=-\sqrt{x}$. Finally, by linearity of expectation, this inequality implies that $E(A-\sqrt{Bg\gamma})\geq 0$, which completes the argument.
Theorem 1.3. Let $\gamma$ be any positive integer; $g_{0}$ be any positive multiple of $\gamma$; and $n$ any multiple of $\gamma g_{0}$. Next, fix any randomized algorithm in our model. Then, there is an input with $n$ jobs, $g$ of which are good for $g\in\{\gamma g_{0},g_{0}/\gamma\}$; and an estimator with estimation gap $\gamma$ on that input. Additionally, the randomized algorithm on that input has expected cost:
@@
Lemma 17 proves that the minimax of this payoff matrix is at least 00. So by Von Neumann’s minimax theorem [98, 127], the maximin is also at least 00. In other words, via Yao’s principal [127], any randomized algorithm has some deterministic input for which $E(A-\sqrt{\gamma Bg})\geq 0$. By linearity of expectation:
@@
Since the sample space of our adversarial distribution has $g\in\{\gamma g_{0},g_{0}/\gamma\}$, the worst-case distribution chosen for the maximin will also have $g\in\{\gamma g_{0},g_{0}/\gamma\}$.
Appendix 0.D Simulation Results
We perform experiments for input distributions where there is a large attack and all good jobs are serviced as late as possible; here, our aim is to verify the scaling behavior predicted by theory. Our custom simulations are written in Python and the source code is available to the public via GitHub [23].
0.D.1 Experiments for LINEAR
@@
Specific Experiments. We conduct two experiments. For our first experiment, we let $n=10\times 2^{x}$ for integer $x\in[-1,18]$. Then, we fix $\gamma=1$ and investigate the impact of different values of $\alpha$ (recall Section 0.A); specifically, we let $\alpha=0,1/2,1$ and $2$.
For our second experiment, we set $n=10\times 2^{x}$, where $x\in[3,18]$. We fix $\alpha=1$ and investigate the impact of different values of $\gamma$; specifically, we let $\gamma=1,2,4,8,16$ and $32$. The range for $x$ is slightly smaller than in the first experiments, since the smallest power-of-$2$-sized input that holds both good and bad jobs is $8010\times 2^{3}=$ 80 when $\gamma=32$.
@@
Results for LINEAR. The results of our first experiment are plotted in Figure 7 (Left). The algorithm’s cost is denoted by $A$, while the adversary’s cost is denoted by $B$. The results show scaling consistent with our theoretical analysis. Notably, setting $\alpha=1$ achieves the most advantageous cost for algorithm versus the adversary. We observe that for $\alpha=0,2$ and $0.5$, the algorithm cost increases by up to a factor of approximately $33.3$, $3.3$, and $1.8$, respectively, relative to LINEAR (i.e., $\alpha=1$) at $B=10^{4}$. We include the line $A=B$ as a reference point, and this closely corresponds to the case for $\alpha=0$ (i.e., each job has cost $1$), since the service cost incurred by the algorithm is close to the number of bad jobs in these experiments. We also include the equation $A=2(\sqrt{B(g+1)}+g)$, which closely fits the line for $\alpha=1$.
@@
0.D.2 Experiments for LINEAR-POWER
@@
For each integer $z\in[10,28]$, we fix a number of bad jobs $b=2^{z}-1$, and let $M$ range over values $1,2,4,$ and $8$. Once all $b$ bad jobs have been serviced—which occurs in period $z$— a final batch of $M$ good jobs arrive with a $1$ hard solution and these are bounced. However, the other $(z+1)M$ (good) jobs will start being serviced. Specifically, $\gamma$ of these (good) jobs will be serviced, with the first $\gamma-1$ jobs (erroneously) considered bad by the estimator, and $\gamma$-th job (correctly) considered good; this ends the first iteration. All remaining good jobs will be serviced in subsequent iterations in similar fashion ($\gamma$ per iteration), since the current price never again exceeds $2^{\left\lfloor\log(\gamma(\gamma+1))\right\rfloor}\leq\gamma(\gamma+1)$ (recall Lemma 2).
Results for LINEAR-POWER. Figure 8 illustrates the results of our simulations when $\gamma=8$. As predicted by our upper bound on cost in Theorem 1.2, LINEAR-POWER achieves a lower cost than the adversary for $B$ sufficiently large. However, aligning with our analysis, we see that this cost advantage decreases as $M$ grows. In particular, between $M=1$ and $M=8$, we see that the cost ratio decreases from roughly $2\times 10^{3}$ to $20$ for a fixed $B=9\times 10^{11}$. Very similar results are observed for other values of $\gamma$, and so we omit these and note that, for this set of experiments, the algorithm’s cost appears to be primarily sensitive to $M$.

### v5

@@ 4 Analysis of LINEAR
By the algorithm, $\hat{g}({\mathcal{I}_{j}})\geq 1$ for all $j$. Let $\mathcal{I}_{tot}=\cup_{1\leq j\leq\ell}\mathcal{I}_{j}$. Using the additive property of $\hat{g}$, we get $\hat{g}(\mathcal{I}_{tot})=\sum_{1\leq j\leq\ell}\hat{g}(\mathcal{I}_{j})\geq\ell$. But by our definition of $\gamma$, we know that $\hat{g}(\mathcal{I}_{tot})\leq\gamma(g(\mathcal{I}_{tot})+1)=\gamma(g+1)$. Thus, $\ell\leq\gamma(g+1)$. ∎
Lemma 2.
@@
Putting together equations 1 and 2 we get that $g(\mathcal{I})\leq\gamma+\gamma^{2}+1$. ∎
Lemma 3.
@@
which proves the lemma. ∎
@@
To bound the adversarial cost in an iteration, we use an integral lower bound. Namely, for any $x$ and $\alpha$, $\sum_{i=1}^{x}i^{\alpha}\geq 1+\int_{i=1}^{x}i^{\alpha}\,di\geq\frac{x^{\alpha+1}}{\alpha+1}$, where the last inequality holds for $\alpha\geq 0$. If the number of iterations is $\ell>0$, we let $x=\left\lfloor b/\ell\right\rfloor$ in the above and bound the total cost of the adversary as:
@@
The second step holds since $(b/\ell)-1=(b-\ell)/\ell$. The fourth step holds since, by Lemma 1, $\ell\leq\gamma(g+1)$. ∎
@@
Let $f:\mathbb{R}\rightarrow\mathbb{R}$ such that, $f(x)=\sum_{j=1}^{\beta}(x+j)^{\alpha}$. For any $i\in[1,\ell]$, this function is continuous when $x\in[b_{i},b_{i}+1]$ and differentiable when $x\in(b_{i},b_{i}+1)$, since it is the sum of a finite number of continuous and differentiable functions respectively. For any $i\in[1,\ell]$, by applying the Mean Value Theorem for the interval $b_{i},b_{i}+1]$, we have:
$\delta(b_{i})$ $=\frac{f(b_{i}+1)-f(b_{i})}{(b_{i}+1)-b_{i}}=f^{\prime}(x_{i})\mbox{~for some~}x_{i}\in[b_{i},b_{i}+1].$
@@
In the above, the last line follows by Lemma 3, and noting that $1/\alpha$ is a fixed constant. ∎
@@
which completes the proof. ∎
@@
Proof.
Since there is no latency, each client immediately receives the current value of PRICE if its first message to the server has a fee with value too small. Then, on the second attempt the client attaches a fee with the correct value, and so receives service. Hence, each good job is serviced with at most $2$ messages from client to server and at most $1$ message from server to client. ∎
Theorem 1 now follows directly from Lemmas 6 and 7.
5 Analysis of LINEAR-POWER
@@
Putting these together, we get that when $B_{i}=0$, $x\leq\log 3\gamma^{2}$ and when $B_{i}>0$, $x\leq((1/2)\log B_{i})+1+\log 3\gamma^{2}$. Hence the maximum current price is at most $3\gamma^{2}\max(1,2\sqrt{B_{i}})\leq 6\gamma^{2}\sqrt{B_{i}+1}$. ∎
Lemma 9.
@@
By Lemma 8, $2^{x}\leq 6\gamma^{2}\sqrt{B_{i}+1}$; taking the log and simplifying, we get $x\leq 3+2\log(\gamma(B_{i}+1))$. The lemma follows since over a period of $2\Delta$ seconds, at most $2M$ good jobs can arrive. ∎
Lemma 10.
@@
If neither the current epoch nor the preceding epoch spanned the beginning of an iteration, then the current price never resets to $1$ during the lifetime of any of these jobs. So the current price is non-decreasing during each such job’s lifetime. Thus, none of the jobs can ever have an attached fee that is larger than the current price, and so no overpayment occurs. ∎
Lemma 11.
@@
By Lemma 8, the maximum current price in either epoch $i$ or i1i-1 is:
@@
This is the maximum amount that any job serviced in epoch $i$ can overpay. Multiplying by the number of good jobs serviced, we obtain the bound in the lemma statement. ∎
Lemma 12.
@@
Lemmas 1314, and 15 below complete the proof of our second main result, Theorem 2.
Lemma 13.
@@
Proof.
Fix the order in which jobs are serviced by LINEAR-POWER. The current price when a job is serviced in LINEAR-POWER will differ by at most a factor of $2$ from LINEAR because of the fact that LINEAR-POWER uses powers of $2$. Thus, the total amount spent by the adversary decreases by at most a factor of $2$, and the total amount spent by the algorithm on all jobs that do not overpay increases by at most a factor of $2$. It follows that the asymptotic results in Theorem 1 hold, except for the costs due to overpayment by the good jobs.
So, we can bound the asymptotic cost of the algorithm as the cost from LINEAR plus the cost due to overpayment. The cost in the theorem statement is the asymptotic sum of the costs from Theorem 1 and from Lemma 12. This is:
@@
which completes the argument. ∎
Lemma 14.
@@
Every time the job is bounced, the latency increases by at most $2\Delta$ seconds: $\Delta$ seconds to send the job to the server, and $\Delta$ seconds to receive the new current price from the server. When the job is serviced it takes at most $\Delta$ seconds: the time to send the job to the server. Thus, the total time for the job to be serviced is at most $2\Delta(2\log(\gamma(B+1))+2)+\Delta\leq\Delta(4\log(\gamma(B+1))+4)$. ∎
@@
Every time a client sends a new message to the server, the fee attached to that message doubles. The maximum current price set by the server over all iterations is at most $\log(B+3\gamma^{2})$, so a good job will send at most that number of messages. Hence, all $g$ good jobs send a total of $O(g\log(B+3\gamma^{2}))$ messages to the server. ∎
6 Lower Bound Proof
@@ 6.2 Analysis
[…]
In the above, the first line follows from Equation 6, and the last line follows from the fact that $\gamma\geq 1$, and the definition of the estimator. Hence, by the above analysis, our estimator has $\gamma$ estimation gap for both distributions. ∎
@@
To show $E(A-\sqrt{Bg\gamma})\geq 0$, we will show that $E(A)/\sqrt{E(B)g}\geq\sqrt{\gamma}$, and then cross-multiply and use Jensen’s inequality. If we define a random variable $a_{i}$ that has value $c_{i}$ if the $i$-th job is good and 0 otherwise, and note that $A=\sum_{i=1}^{n}a_{i}$, then by linearity of expectation, the expected algorithm cost due to fees in distribution $1$ is $\sum_{i=1}^{n}E(a_{i})=C\gamma g_{0}/n$, with service cost of $n$. A similar analysis for $E_{1}(B)$ gives:
@@
where the second line holds by applying Jensen’s inequality [87] to the convex function [^2] Equivalently, we can directly apply the (less well-known) version of Jensen’s inequality for concave functions [43] to obtain the second line. $f(x)=-\sqrt{x}$. Finally, by linearity of expectation, this inequality implies that $E(A-\sqrt{Bg\gamma})\geq 0$, which completes the argument. ∎
Theorem 3. Let $\gamma$ be any positive integer; $g_{0}$ be any positive multiple of $\gamma$; and $n$ any multiple of $\gamma g_{0}$. Next, fix any randomized algorithm in our model. Then, there is an input with $n$ jobs, $g$ of which are good for $g\in\{\gamma g_{0},g_{0}/\gamma\}$; and an estimator with estimation gap $\gamma$ on that input. Additionally, the randomized algorithm on that input has expected cost:
@@
Lemma 17 proves that the minimax of this payoff matrix is at least 0. So by Von Neumann’s minimax theorem [95, 124], the maximin is also at least 0. In other words, via Yao’s principal [124], any randomized algorithm has some deterministic input for which $E(A-\sqrt{\gamma Bg})\geq 0$. By linearity of expectation:
@@
Since the sample space of our adversarial distribution has $g\in\{\gamma g_{0},g_{0}/\gamma\}$, the worst-case distribution chosen for the maximin will also have $g\in\{\gamma g_{0},g_{0}/\gamma\}$. ∎
7 Preliminary Simulation Results
We conduct simulations for input distributions where there is a large attack and all good jobs are serviced as late as possible; here, our aim is to simply verify the scaling behavior predicted by theory. Our custom simulations are written in Python and the source code is available to the public via GitHub [23].
7.1 Experiments for LINEAR
@@
Specific Experiments. We conduct two experiments. For our first experiment, we let $n=10\times 2^{x}$ for integer $x\in[-1,18]$. Then, we fix $\gamma=1$ and investigate the impact of different values of $\alpha$ (recall Section 4); specifically, we let $\alpha=0,1/2,1$ and $2$.
For our second experiment, we set $n=10\times 2^{x}$, where $x\in[3,18]$. We fix $\alpha=1$ and investigate the impact of different values of $\gamma$; specifically, we let $\gamma=1,2,4,8,16$ and $32$. The range for $x$ is slightly smaller than in the first experiments, since the smallest power-of-$2$-sized input that holds both good and bad jobs is $10\times 2^{3}=80$ when $\gamma=32$.
@@
Results for LINEAR. The results of our first experiment are plotted in Figure 7 (Left). The algorithm’s cost is denoted by $A$, while the adversary’s cost is denoted by $B$. The results show scaling consistent with our theoretical analysis. Notably, setting $\alpha=1$ achieves the most advantageous cost for algorithm versus the adversary. We observe that for $\alpha=0,2$ and $0.5$, the algorithm cost increases by up to a factor of approximately $33.3$, $3.3$, and $1.8$, respectively, relative to LINEAR (i.e., $\alpha=1$) at $B=10^{4}$. We include the line $A=B$ as a reference, and this closely corresponds to the case for $\alpha=0$ (i.e., each job has cost $1$), since the service cost incurred by the algorithm is close to the number of bad jobs in these experiments. We also include the equation $A=2(\sqrt{B(g+1)}+g)$, which closely fits the line for $\alpha=1$.
@@
7.2 Experiments for LINEAR-POWER
@@
For each integer $z\in[10,28]$, we fix a number of bad jobs $b=2^{z}-1$, and let $M$ range over values $1,2,4,$ and $8$. Once all $b$ bad jobs have been serviced—which occurs in period zz— a final batch of $M$ good jobs arrive with a $1$ hard solution and these are bounced. However, the other $(z+1)M$ (good) jobs will start being serviced. Specifically, $\gamma$ of these (good) jobs will be serviced, with the first $\gamma-1$ jobs (erroneously) considered bad by the estimator, and $\gamma$-th job (correctly) considered good; this ends the first iteration. All remaining good jobs will be serviced in subsequent iterations in similar fashion ($\gamma$ per iteration), since the current price never again exceeds $2^{\left\lfloor\log(\gamma(\gamma+1))\right\rfloor}\leq\gamma(\gamma+1)$ (recall Lemma 2).
Results for LINEAR-POWER. Figure 8 illustrates the results of our simulations when $\gamma=8$. As predicted by our upper bound on cost in Theorem 2, LINEAR-POWER achieves a lower cost than the adversary for $B$ sufficiently large. However, aligning with our analysis, we see that this cost advantage decreases as $M$ grows. In particular, between $M=1$ and $M=8$, we see that the cost ratio decreases from roughly $2\times 10^{3}$ to $20$ for a fixed $B=9\times 10^{11}$. Very similar results are observed for other values of $\gamma$, and so we omit these and note that, for this set of experiments, the algorithm’s cost appears to be primarily sensitive to $M$.
@@
8 Conclusion and Future Work
Many interesting problems remain. Are there less restrictive properties for the estimator which suffice to achieve a good pricing strategy? Can we model both the clients and the server as selfish but rational agents? In particular, can we design a mechanism that results in a Nash equilibrium, and also provides good cost performance as a function of the adversarial cost?
Acknowledgements. This material is based upon work supported by the National Science Foundation under grant numbers CNS-2210299, CNS-2210300, and CCF-2144410.